# Bijective proof of a conjecture on unit interval posets

### 12 Jan 2024 v2

@@ 1 Introduction
The study of unit interval posets started in statistics (see [WF57]) and psychology (see [Sco64]). However, it has since been connected to other objects in algebraic combinatorics, such as $(3+1)$-free posets [SR03] and positroids [CG18]. The connection to $(3+1)$-free posets is particularly important, as these posets are at the center of the Stanley–Stembridge conjecture [SS93], which states that the chromatic symmetric function of the incomparability graph of a $(3+1)$-free poset has only positive coefficients when expanded in the basis of elementary symmetric functions. Unit interval posets thus receive some attention as it can be used to study the structure of $(3+1)$-free posets [SR03, GPMR14, LZ13].
It is known that unit interval posets are counted by Catalan numbers [WF57], and researchers have given two different bijections to represent a unit interval poset by Dyck paths [SR03, GPMR14]. In a recent preprint of Matherne, Morales and Selover [MMS], it was conjectured that the two bijections are related by the famous zeta map in $q,t$-Catalan combinatorics (see [Hag08]), which was first given in [AKOP02]. In this short note, we settle this conjecture using bijections (see 4.6), in contrast to a recent inductive proof of the same conjecture by Gélinas, Segovia and Thomas [GST]. To this end, we first introduce a bijection $\Xi_{\mathrm{poset}}$ between unit interval posets and plane trees. We then show that the two different known bijections from unit interval posets are conjugates of special cases of some bijections $\Xi_{\mathrm{steep}}$ and $\Xi_{\mathrm{bounce}}$ in [CFM20] constructed for the study of parabolic Tamari lattices. Using the link between $\Xi_{\mathrm{steep}},\Xi_{\mathrm{bounce}}$ and the zeta map also established in the same article, we conclude our proof.
The rest of the article is organized as follows. In Section 2 we give the basic definitions. Then we give our bijection between unit interval posets and plane trees in Section 3, along with the related bijections in [CFM20]. Finally in Section 4 we recall the two known bijections between unit interval posets and Dyck paths, and then establish their link with the bijections in Section 3, leading to a bijective proof of our main result (4.6).
@@ 2 Preliminary
Theorem 2.1 (Theorem 2.1 of [Sco64]).
A poset $P$ is a unit interval poset if and only if it is $(3+1)$-free and $(2+2)$-free, that is, the order induced on any four elements cannot be a chain of $3$ elements plus an incomparable elements, or two disjoint chains each containing $2$ elements.
@@ 3 Unit interval posets and plane trees
Suppose that $T$ has maximal arity $m$. Let $v^{\prime}$ be the parent of $u$. By (1), we have $1<x_{u}-x_{v^{\prime}}<1+(m+2)^{-\ell+1}$, with $\ell$ the distance from $v^{\prime}$ to the root. Then, $x_{v}$ satisfying $x_{v^{\prime}}<x_{v}<1+x_{u}$ means $0<x_{v}-x_{v^{\prime}}<(m+2)^{-\ell+1}$, which means $v=v^{\prime}$ by (1), as it requires $v$ and $v^{\prime}$ to have the same distance to the root, but in this case $|x_{v}-x_{v^{\prime}}|$ cannot be smaller than $(m+2)^{-\ell+1}$ if $v\neq v^{\prime}$. We thus have the equivalence. ∎
@@
Let $n\geq 1$ and $T\in\mathcal{T}_{n}$, we construct a Dyck path $D$ by appointing the number of north steps at each $x$-coordinate (see the Dyck path on the left of Figure 2). Let $\alpha_{\ell}$ be the number of nodes of depth $\ell$ (thus distance $\ell$ to the root), and $\ell_{\max}$ the maximal depth of nodes in $T$. We take $s_{\ell}=\alpha_{1}+\cdots+\alpha_{\ell}$. Given $11\leq k\leq n-$ 1, there is a unique way to write $k=s_{\ell}-r$ with $0\leq r<\alpha_{\ell}$ and $1\leq\ell\leq\ell_{\max}$. In this case, the number of north steps in $D$ on $x=k$ is the number of children of the $(r+1)$-st node of depth $\ell$. The number of north steps in $D$ on $x=0$ is the number of children of the root. We define $\Xi_{\mathrm{bounce}}(T)=D$. The map $\Xi_{\mathrm{bounce}}$ is a bijection from $\mathcal{T}_{n}$ to $\mathcal{D}_{n}$ for all $n\geq 1$ (see Lemma 3.18 in [CFM20]), and we denote its inverse by $\Lambda_{\mathrm{bounce}}=\Xi_{\mathrm{bounce}}^{-1}$.
Proposition 3.7 (Theorem IV of [CFM20]).
@@
See Figure 2 for an illustration of 3.7. We do not give the original definition of the zeta map $\zeta$ in [AKOP02] here, and will instead be using 3.7 in the following, as it is better suited to our approach.
@@
Albeit the same notation, the maps $\Xi_{\mathrm{steep}}$ and $\Xi_{\mathrm{bounce}}$ are only special cases of the ones in [CFM20, Construction 3.10 and Equation 14]. More precisely, the central objects in [CFM20] are called LAC trees, which are plane trees with non-root nodes divided into regions algorithmically. We are only using the special case where the region $d$ consists of nodes of depth $d$, which is also the one used in [CFM20, Section 3.3] to prove 3.7. In full generality, the map $\Xi_{\mathrm{steep}}$ maps a LAC tree to a steep pair, which is a pair of nested Dyck paths with a steep upper path, i.e.,without two consecutive east steps except on the maximal height. The map $\Xi_{\mathrm{bounce}}$ maps a LAC tree to a bounce pair, which is a pair of nested Dyck path with a bounce lower path, i.e., a concatenation of sub-paths of the form $\uparrow^{k}\rightarrow^{k}$ for all $k$. This explains the words “steep” and “bounce” in their notation.
@@
The bijection $\Xi_{\mathrm{steep}}$ is classical, except that we do the contour walk from right to left instead of from left to right. The bijection $\Xi_{\mathrm{bounce}}$ is quite close to the classical bijection between plane trees and Łukasiewicz words (see [FS09, Section I.5.3]), which can be seen as sequences $(z_{0},z_{1},\ldots,z_{n-1})$ with $z_{i}\geq-1$ such that $\sum_{i=0}^{k}z_{i}\geq 0$ for $10\leq k<n-$ 1, and the sum of all $z_{i}$’s is $-1$. However, in the classical bijection of Łukasiewicz, we deal with nodes in a depth-first order, but in $\Xi_{\mathrm{bounce}}$ they are processed in a breadth-first order.
@@ 4 Unit interval posets and the zeta map
We start by the first representation, which was first defined implicitly using anti-adjacency matrices of unit interval posets in [SR03], only involving the poset structure. The following form was first given in [CG18].
@@
Although defined here using the set $S$, the map $\varphi$ in 4.1, with this form given in [CG18, Section 5] and proved in Lemma 5.7 therein to be equivalent to the original definition, which is the one used in [MMS, Section 2.3], does not depend on the choice of $S$.
@@
For the line $x=0$, we observe that the number of elements in $S$ strictly smaller than $x_{1}+1$ must be those with integer part equal to $1$, thus those from the children of the root of $T$. For $11\leq k\leq n-$ 1, from 4.1, we know that the number of north steps on the line $x=k$ is the number of elements $x_{i}$ such that $x_{k-1}+1<x_{i}<x_{k}+1$, as $x_{k}+1$ corresponds to the $k$-th east step in $D$. Let $u_{k}$ be the node in $T$ corresponding to $x_{k}$. By 3.3, the nodes in $T$ corresponding to such $x_{i}$ are children of $u_{k}$, meaning that the total number of north steps on $x=k$ is the number of children of $u_{k}$. Furthermore, suppose that $u_{k}$ is of depth $\ell$, then we can write $k=s_{\ell}-r$ with $0\leq r<\alpha_{\ell}$ as in 3.6. We know that $u_{k}$ is the $(r+1)$-st node of depth $\ell$ in $T$, as in 3.1 the values corresponding to nodes of the same depth are decreasing from left to right. We thus conclude that $D=D^{\prime}$, and $\varphi\circ\Xi_{\mathrm{poset}}=\Xi_{\mathrm{bounce}}$. As both $\Xi_{\mathrm{bounce}}$ ([CFM20, Lemma 3.18] and $\Xi_{\mathrm{poset}}$ (3.4) are bijections, we conclude that $\varphi$ is also a bijection. ∎
The second presentation was first defined in [GPMR14] for $(3+1)$-free posets, and it takes a simpler form for unit interval posets. A more explicit presentation is given in [GP13, Section 2].
@@
The following result ensures that $\psi$ is well-defined. It is a special case of [MMS, Theorem 5.2], which is obtained by combining [GPMR14, Remark 3.2, Proposition 3.11]. A self-contained direct proof can be found in [GST, Section 6].
Proposition 4.4 (Special case of [MMS, Theorem 5.2]).
@@
Theorem 4.6 (Conjecture 7.1 in [MMS]).

### 22 Feb 2024 v3

@@ 1 Introduction
The study of unit interval posets started in statistics (Wine and Freund, 1957) and psychology (Scott, 1964). However, it has since been connected to other objects in algebraic combinatorics, such as $(3+1)$-free posets Skandera and Reed (2003) and positroids Chavez and Gotti (2018). The connection to $(3+1)$-free posets is particularly important, as these posets are at the center of the Stanley–Stembridge conjecture in Stanley and Stembridge (1993), which states that the chromatic symmetric function of the incomparability graph of a $(3+1)$-free poset has only positive coefficients when expanded in the basis of elementary symmetric functions. Unit interval posets thus receive some attention as it can be used to study the structure of $(3+1)$-free posets (Skandera and Reed, 2003; Guay-Paquet, Morales, and Rowland, 2014; Lewis and Zhang, 2013).
It is known in Wine and Freund (1957) that unit interval posets are counted by Catalan numbers, and researchers have given two different bijections to represent a unit interval poset by Dyck paths (Skandera and Reed, 2003; Guay-Paquet, Morales, and Rowland, 2014). In a recent preprint (Matherne, Morales, and Selover, ), it was conjectured that the two bijections are related by the famous zeta map in $q,t$-Catalan combinatorics (Haglund, 2008), which was first given in (Andrews, Krattenthaler, Orsina, and Papi, 2002). In this short note, we settle this conjecture using bijections (see 4.6), in contrast to a recent inductive proof of the same conjecture by Gélinas, Segovia, and Thomas . To this end, we first introduce a bijection $\Xi_{\mathrm{poset}}$ between unit interval posets and plane trees. We then show that the two different known bijections from unit interval posets are conjugates of special cases of some bijections $\Xi_{\mathrm{steep}}$ and $\Xi_{\mathrm{bounce}}$ from Ceballos, Fang, and Mühle (2020) constructed for the study of parabolic Tamari lattices. Using the link between $\Xi_{\mathrm{steep}},\Xi_{\mathrm{bounce}}$ and the zeta map established in the same article, we conclude our proof.
The rest of the article is organized as follows. In Section 2 we give the basic definitions. Then we give our bijection between unit interval posets and plane trees in Section 3, along with the related bijections from Ceballos, Fang, and Mühle (2020). Finally in Section 4 we recall the two known bijections between unit interval posets and Dyck paths, and then establish their link with the bijections in Section 3, leading to a bijective proof of our main result (4.6).
@@ 2 Preliminary
Theorem 2.1 (Scott (1964, Theorem 2.1)).
A poset $P$ is a unit interval poset if and only if it is $(3+1)$-free and $(2+2)$-free, that is, the order induced on any four elements cannot be a chain of $3$ elements plus an incomparable element, or two disjoint chains each containing $2$ elements.
@@ 3 Unit interval posets and plane trees
Suppose that $T$ has maximal arity $m$. Let $v^{\prime}$ be the parent of $u$. By (1), we have $1<x_{u}-x_{v^{\prime}}<1+(m+2)^{-\ell+1}$, with $\ell$ the distance from $v^{\prime}$ to the root. Then, $x_{v}$ satisfying $x_{v^{\prime}}\leq x_{v}<1+x_{u}$ means $0\leq x_{v}-x_{v^{\prime}}<(m+2)^{-\ell+1}$, which means $v=v^{\prime}$ by (1), as it requires $v$ and $v^{\prime}$ to have the same distance to the root, but in this case $|x_{v}-x_{v^{\prime}}|$ cannot be smaller than $(m+2)^{-\ell+1}$ if $v\neq v^{\prime}$. We thus have the equivalence. ∎
@@
Let $n\geq 1$ and $T\in\mathcal{T}_{n}$, we construct a Dyck path $D$ by appointing the number of north steps at each $x$-coordinate (see the Dyck path on the left of Figure 2). Let $\alpha_{\ell}$ be the number of nodes of depth $\ell$ (thus distance $\ell$ to the root), and $\ell_{\max}$ the maximal depth of nodes in $T$. We take $s_{\ell}=\alpha_{1}+\cdots+\alpha_{\ell}$. Given $11\leq k\leq n-$ 1, there is a unique way to write $k=s_{\ell}-r$ with $0\leq r<\alpha_{\ell}$ and $1\leq\ell\leq\ell_{\max}$. In this case, the number of north steps in $D$ on $x=k$ is the number of children of the $(r+1)$-st node of depth $\ell$. The number of north steps in $D$ on $x=0$ is the number of children of the root. We define $\Xi_{\mathrm{bounce}}(T)=D$. The map $\Xi_{\mathrm{bounce}}$ is a bijection from $\mathcal{T}_{n}$ to $\mathcal{D}_{n}$ for all $n\geq 1$, see Ceballos, Fang, and Mühle (2020, Lemma 3.18), and we denote its inverse by $\Lambda_{\mathrm{bounce}}=\Xi_{\mathrm{bounce}}^{-1}$.
Proposition 3.7 (Ceballos, Fang, and Mühle (2020, Theorem IV)).
@@
See Figure 2 for an illustration of 3.7. Here, we do not give the original definition of the zeta map $\zeta$ in Andrews, Krattenthaler, Orsina, and Papi (2002), and will instead be using 3.7 in the following, as it is better suited to our approach.
@@
Albeit the same notation, the maps $\Xi_{\mathrm{steep}}$ and $\Xi_{\mathrm{bounce}}$ are only special cases of the ones from Ceballos, Fang, and Mühle (2020, Construction 3.10 and Equation 14). More precisely, the central objects of Ceballos, Fang, and Mühle (2020) are called LAC trees, which are plane trees with non-root nodes divided into regions algorithmically. We are only using the special case where the region $d$ consists of nodes of depth $d$, which is also the one used in Ceballos, Fang, and Mühle (2020, Section 3.3) to prove 3.7. In full generality, the map $\Xi_{\mathrm{steep}}$ maps a LAC tree to a steep pair, which is a pair of nested Dyck paths with a steep upper path, i.e.,without two consecutive east steps except on the maximal height. The map $\Xi_{\mathrm{bounce}}$ sends a LAC tree to a bounce pair, which is a pair of nested Dyck path with a bounce lower path, i.e., a concatenation of sub-paths of the form $\uparrow^{k}\rightarrow^{k}$ for all $k$. This explains the words “steep” and “bounce” in their notation.
@@
The bijection $\Xi_{\mathrm{steep}}$ is classical, except that we do the contour walk from right to left instead of from left to right. The bijection $\Xi_{\mathrm{bounce}}$ is quite close to the classical bijection between plane trees and Łukasiewicz words, as given in Flajolet and Sedgewick (2009, Section I.5.3)), which can be seen as sequences $(z_{0},z_{1},\ldots,z_{n-1})$ with $z_{i}\geq-1$ such that $\sum_{i=0}^{k}z_{i}\geq 0$ for $10\leq k<n-$ 1, and the sum of all $z_{i}$’s is $-1$. However, in the classical bijection of Łukasiewicz, we deal with nodes in a depth-first order, but in $\Xi_{\mathrm{bounce}}$ they are processed in a breadth-first order.
@@ 4 Unit interval posets and the zeta map
We start by the first representation, which was first defined implicitly using anti-adjacency matrices of unit interval posets in Skandera and Reed (2003), only involving the poset structure. The following form was first given in Chavez and Gotti (2018).
@@
This form of the map $\varphi$ in 4.1 was given in Chavez and Gotti (2018, Section 5) and proved in Lemma 5.7 therein to be equivalent to the original definition, which is the one used in Matherne, Morales, and Selover (, Section 2.3). Although defined here using the set $S$, the map $\varphi$ does not depend on the choice of $S$.
@@
For the line $x=0$, we observe that the number of elements in $S$ strictly smaller than $x_{1}+1$ must be those with integer part equal to $1$, thus those from the children of the root of $T$. For $11\leq k\leq n-$ 1, from 4.1, we know that the number of north steps on the line $x=k$ is the number of elements $x_{i}$ such that $x_{k-1}+1<x_{i}<x_{k}+1$, as $x_{k}+1$ corresponds to the $k$-th east step in $D$. Let $u_{k}$ be the node in $T$ corresponding to $x_{k}$. By 3.3, the nodes in $T$ corresponding to such $x_{i}$ are children of $u_{k}$, meaning that the total number of north steps on $x=k$ is the number of children of $u_{k}$. Furthermore, suppose that $u_{k}$ is of depth $\ell$, then we can write $k=s_{\ell}-r$ with $0\leq r<\alpha_{\ell}$ as in 3.6. We know that $u_{k}$ is the $(r+1)$-st node of depth $\ell$ in $T$, as in 3.1 the values corresponding to nodes of the same depth are decreasing from left to right. We thus conclude that $D=D^{\prime}$, and $\varphi\circ\Xi_{\mathrm{poset}}=\Xi_{\mathrm{bounce}}$. As both $\Xi_{\mathrm{bounce}}$ (Ceballos, Fang, and Mühle, 2020, Lemma 3.18) and $\Xi_{\mathrm{poset}}$ (3.4) are bijections, we conclude that $\varphi$ is also a bijection. ∎
The second presentation was first defined in Guay-Paquet, Morales, and Rowland (2014) for $(3+1)$-free posets, and it takes a simpler form for unit interval posets. A more explicit presentation is given in Guay-Paquet (2013, Section 2).
@@
The following result ensures that $\psi$ is well-defined. It is a special case of Matherne, Morales, and Selover (, Theorem 5.2), which is obtained by combining Guay-Paquet, Morales, and Rowland (2014, Remark 3.2, Proposition 3.11). A self-contained direct proof of this special case can be found in Gélinas, Segovia, and Thomas (, Section 6).
Proposition 4.4 (Special case of Matherne, Morales, and Selover (, Theorem 5.2)).
@@
Theorem 4.6 (Matherne, Morales, and Selover (, Conjecture 7.1)).